# Leveraging Moving Sound Source Trajectories for Universal Sound Separation

### v1

@@ Abstract
Existing methods utilizing spatial information for sound source separation require prior knowledge of the direction of arrival (DOA) of the source or utilize estimated but imprecise localization results, which impairs the separation performance, especially when the sound sources are moving. In fact, sound source localization and separation are interconnected problems, that is, sound source localization facilitates sound separation while sound separation contributes to more precise source localization. This paper proposes a method utilizing the mutual facilitation mechanism between sound source localization and separation for moving sources. Initially, sound separation is conducted using rough preliminary sound source tracking results. Sound source tracking is then performed on the separated signals thus the tracking results can become more precise. The precise trajectory can further enhances the separation performance. This mutual facilitation process can be performed over several iterations. Simulation experiments conducted under reverberation conditions and with moving sound sources demonstrate that the proposed method can achieve more accurate separation based on more precise tracking results.
@@ I Introduction
Humans are capable of accurately localizing a sound source and separating the sound in complex and noisy environments. This phenomenon is referred to as the ”cocktail party effect”. The task of separating the sound of each universal source from the audio mixture is called universal sound separation. In an actual sound separation system, the spatial information of the sound sources especially the localization results can further assist in sound separation. Exploring how to utilize spatial information for separation is a worthwhile research problem. Traditional sound separation algorithms that utilize spatial location information are mostly based on beamforming techniques, which suppress the signals from directions other than the target direction. These methods often perform poorly under reverberant conditions. Additionally, they depend on the accurate estimation of the noise covariance matrix and rely on prior information about the target signal.
@@
Recently, Deep neural networks have shown impressive efficacy in sound separation tasks[1][2][3][4]. Utilizing spatial information in sound separation with neural networks has been investigated. The typical approach involves cascading single-channel separation with beamforming techniques like minimum variance distortionless response (MVDR) beamformers [5][6][7][8]. These methods use the separation network to generate the masks of target sound and interfering sound on each channel, and the spatial covariance matrices are then calculated to generate the beamformer filters. These methods heavily rely on the performance of single-channel separation and spatial information does not effectively aid in single-channel separation; when mask estimation of the target sound is inaccurate, the performance of the beamformer degrades significantly.
@@
Some methods use neural networks to estimate the weights of the filters of beamformers. Luo et al. proposed FasNet to estimate the weights of the spatial filters with multi-channel mixture signals [9]. Subsequently, the Transform-Average-Concatenate (TAC) mechanism is introduced in FasNet to avoid selecting a reference microphone [10]. However, the performance of this method degrades under reverberant conditions. Besides, the location information of the sound source has not been utilized to assist in the separation. In [11], an all deep learning MVDR framework is proposed to estimate every component in the calculation of a MVDR filter. [12] utilizes the inter-channel phase difference (IPD) to convert the DOA into direction features to direct the separation network. This kind of method requires prior knowledge of the DOA of the sound source, and its performance significantly degrades when there is a large error in DOA estimation. Some neural beamformer methods don’t need to know the sources’ DOA, but they require the sources to remain stationary [13][14].
@@
Another kind of method to utilize the spatial information is training the multi-channel separation networks end to end with feature extraction across the channel dimension. [15] extends the Conv-Tasnet [4] and extracts the inter-channel features with 2D-TCN layers. A triple-path method utilizing Squeeze-and-Excitation mechanism [16] across the channel dimension to learn the spatial feature is introduced in [17]. In [18], the dual-path transformer method [19] is further extended to a triple-path attentive recurrent network in which the ARN module consisting of a RNN block, an attention block and a feedforword block is applied on the inter-channel, intra-chunk and inter-chunk dimension. TF-GridNet is proposed using 2D-convolutional layers to extract spatial information, bi-directional Long Short-Term Memory networks (BLSTM) to extract intra-frame and sub-band information and multi-head self-attention to extract full-band information in [20]. SpatialNet is proposed in [21] which utilizes the convolutional layers to extract spatial information followed by cross-band blocks and narrow-band blocks operated on T-F dimension and outperforms TF-GridNet with less parameters. The state-space model [22] is further introduced to enhance training and inference speed in [23]. However, such methods only exploiting spatial information implicitly by leveraging the multi-channel information and do not effectively utilize the DOA information of the sound source. Besides, these methods assume that the sound sources hardly move.
@@
In fact, humans can effectively utilize spatial information in complex environments to simultaneously localize and separate sound sources. The localization results can aid in sound source separation, and the separation results can also assist the localization. Localization and separation are mutually supportive processes. Some studies have explored how to leverage both localization and separation simultaneously. In [24], the spatial area is initially divided into several regions, the sound of sources in each region is separated, and the boundaries of these regions are iteratively refined through binary search based on the iterative separation results. However, this method relies on the assumption that the sound source is stationary. Hassan Taherian et al. performed localization based on multi-channel separation and used the localization results to determine the number of sound sources in the current frame, making the subsequent continuous separation more precise [25]. But this method does not incorporate the localization results into the separation process. LaSNet was proposed in [26], which integrates localization and separation within a single network, using loss functions for both tasks to jointly optimize the network. This approach considers localization and separation as a multi-task learning problem, leveraging only the shared information between the two tasks in the hidden layer space, without clearly defining the mechanism by which the two tasks assist each other. This method also requires the sources to remain stationary.
@@
Most of the sound separation methods assume the sound sources to hardly move. Approaches for handling moving sound sources mostly rely on dynamically estimating the location of the sound source. Some methods estimate the location of the sources with the steered response power (SRP) methods on each frame first and track the sources based on the estimated locations. The dynamic locations is utilized as steer vectors and the sound is separated based on them [27][28]. The localization results may deteriorate under reverberant environments, which can significantly impact the separation performance. Besides, these methods does not leverage the relevant information between localization and separation. Hokuto et al. jointly perform the sound separation and localization with an inference model with time-variant spatial covariance matrices (SCMs) and introduce the temporal smoothness of SCMs to the inference model [29]. But this method relies on assumptions about the prior distribution of the data and the sound source separation and localization don’t benefit from each other.
In this paper, a method utilizing Mutual Facilitation mechanism between sound source Tracking and Separation (MFTS) is proposed. Considering that when the energy of a sound source is low, its localization accuracy significantly decreases, making source separation based on such localization results meaningless, we first estimate the envelopes of the energy of the sound sources, which can be viewed as simplified sound separation. These estimated envelopes then assist the sound source tracking. The initial tracking trajectory can serve as the starting condition for mutual facilitation. During the mutual facilitation, the separated multi-channel signal only contains one source can be obtained based on the estimated trajectory. This separation results can further yield a more precise trajectory, which in turn can further enhance the separation results.
@@
The rest of this paper is organized as follows. Section II introduces the problem formulation of the target sound extraction. Section III presents the proposed method, including the tracking and sound extraction modules, and shows how they collaborate together. Section IV verifies the effectiveness of the method through simulation experiments. Finally, our conclusion is presented in section V.
@@ II Universal Sound Separation
$\begin{split}y_{m}(n)&=\sum_{c=1}^{C}x_{cm}(n)\\ &=\sum_{c=1}^{C}s_{c}(n)*h_{cm}(n),\end{split}$ (1)
where $y_{m}(n)$ is the mixture signal of microphone $m$ at time $n$, $C$ is the number of the sources, $x_{cm}$ is the signal at the microphone $m$ from the $c-th$ source, $s_{c}(n)$ is the signal of source $c$ and $h_{cm}(n)$ is the room impulse responses (RIR) between source $c$ and microphone $m$. The goal of universal sound separation is to separate the signal $x_{cm}$ of sound source $c$ at reference channel $m$. The location of the sound sources is a condition that can effectively aid in sound separation. By providing the location or DOA of source $c$ as additional input, the sound of the target source $c$ at microphone $m$ can be extracted.
The choice of reference microphone can also affect the performance of sound source separation. To eliminate the impact of reference microphone selection on separation performance for sources at different locations, we use First-Order Ambisonics (FOA) signals and the zeroth-order signal is selected as the separation target. The zeroth-order ambisonic signal is equivalent to the signal recorded by an omnidirectional microphone placed at the center of a spherical microphone array, thereby neutralizing the performance variations due to different source positions and reference microphone selections.
Figure 1: The structure of proposed method. (a) Stage 1: the initial tracking module. (b) Stage 2: the mutual facilitation module. (c) Stage 3: the neural beamforming module.
@@ III Method
The structure of our proposed method is shown in Fig 1. The proposed method can be divided into three stages. In the first stage, we perform initial tracking of all the sound sources based on the estimated envelope of each source’s energy. These initial trajectories might not be highly precise. In the second stage, a mutual facilitation process for source tracking and target sound extraction is performed. Specifically, the initial trajectory is used as a condition and input together with the mixture signals into the target source extraction network to extract the multi-channel sound corresponding to the trajectory. The separation result is subsequently input to precise tracking network which only tracks one source based on the multi-channel signal. After that, this precise trajectory is then used as a condition to derive more precise multi-channel separation results based on which a more precise trajectory can be then derived. In the third stage, a neural-beamformer utilizes the precise separation and tracking results to produce the single-channel sound separation signal. The training mechanisms for the above networks are described as follows.
@@ III-A Initial tracking
The initial tracking module takes mixture signals as input and outputs the trajectories of the sources. We first estimate the envelope of the energy of each sound source in the mixture. This envelope is then input to the tracking network along with the audio mixture signals. Concurrently, the tracking network’s output is multiplied by the envelope to compute the loss.
@@
The envelope of the source energy is generated by a maximum filter calculating the maximum of the absolute value of the audio signal in each frame, which can be denoted as $E_{m,trg}\in R^{1\times N}$, where $N$ is the number of frames. The zero-th order of FOA signal is used to generate the envelope. The envelope of the sound is obtained by linearly interpolating $E_{m}^{trg}$, which is denoted as $E^{trg}\in R^{1\times T}$, where $T$ is the number of temporal samples of input audio signal. $T=samplerate\times seconds$. The network used for estimating the envelope in this study is SpatialNet [21], which is a network developed for sound separation. SpatialNet accepts multi-channel mixed signals as input and outputs single-channel separated signal for all sound sources. We select this network as the task of estimating the energy envelope of each sound source from the audio mixture can be considered a simplified version of sound separation. SpatialNet employs 1D convolutions to extract spatial and crossband information and uses a multi-head attention mechanism along with time convolutional layers (T-Conv) to capture narrowband information. We follow the approach in [23] by replacing the attention and T-Conv modules with the recently proposed Mamba [22] to enhance computational efficiency. We add a max-pooling layer at the end of the network, applied to the absolute value of the original network output, to obtain the maximum value points for each frame consistent with the labels. The loss function for training the envelope estimation network is Normalized Mean Square Error (NMSE), which is denoted as:
$\begin{split}Loss_{env}&=NMSELoss(E_{m}^{est},E_{m}^{trg})\\ &=10log_{10}\frac{\sum_{n=0}^{N}(E_{m}^{est}[n]-E_{m}^{trg}[n])^{2}}{\sum_{n=0}^{N}(E_{m}^{trg}[n])^{2}},\end{split}$ (2)
@@
where $E_{m}^{est}$ is the estimated maximum points for calculating the envelope. Besides, utterance-level permutation invariant training (uPIT) is applied during training. The estimated envelope is also obtained by linearly interpolating $E_{m}^{est}$, which is denoted as $E^{est}\in R^{1\times T}$.

### v2

# Leveraging Sound Source Trajectories for Universal Sound Separation

@@ Abstract
Existing methods utilizing spatial information for sound source separation require prior knowledge of the direction of arrival (DOA) of the source or utilize estimated but imprecise localization results, which impairs the separation performance, especially when the sound sources are moving. In fact, sound source localization and separation are interconnected problems, that is, sound source localization facilitates sound separation while sound separation contributes to refined source localization. This paper proposes a method utilizing the mutual facilitation mechanism between sound source localization and separation for moving sources. The proposed method comprises three stages. The first stage is initial tracking, which tracks each sound source from the audio mixture based on the source signal envelope estimation. These tracking results may lack sufficient accuracy. The second stage involves mutual facilitation: Sound separation is conducted using preliminary sound source tracking results. Subsequently, sound source tracking is performed on the separated signals, thereby refining the tracking precision. The refined trajectories further improve separation performance. This mutual facilitation process can be iterated multiple times. In the third stage, a neural beamformer estimates precise single-channel separation results based on the refined tracking trajectories and multi-channel separation outputs. Simulation experiments conducted under reverberant conditions and with moving sound sources demonstrate that the proposed method can achieve more accurate separation based on refined tracking results.
@@ I Introduction
Humans are capable of accurately localizing a sound source and separating the sound in complex and noisy environments. This phenomenon is referred to as the ”cocktail party effect”. The task of separating the sound of each universal source from the audio mixture is called universal sound separation. In an actual sound separation system, the spatial information of the sound sources, especially the localization results, can further assist in sound separation. Exploring how to utilize spatial information for separation is a worthwhile research problem. Traditional sound separation algorithms that utilize spatial location information are mostly based on beamforming techniques, which suppress the signals from directions other than the target direction [1]. These methods often perform poorly under reverberant conditions. Additionally, they depend on the accurate estimation of the noise covariance matrix and rely on prior information about the target signal.
@@
Recently, deep neural networks have shown significant efficacy in sound separation tasks [2],[3],[4],[5]. Utilizing spatial information in sound separation with neural networks has been investigated. The typical approach involves cascading single-channel separation with beamforming techniques like minimum variance distortionless response (MVDR) beamformers [6],[7],[8],[9],[10]. These methods use the separation network to generate the masks of target sound and interfering sound on each channel, and the spatial covariance matrices are then calculated to generate the beamformer filters. These methods heavily rely on the performance of single-channel separation and spatial information does not effectively aid in single-channel separation; when mask estimation of the target sound is inaccurate, the performance of the beamformer degrades significantly.
@@
Some methods use neural networks to estimate the weights of the filters of beamformers. Luo et al. proposed FasNet to estimate the weights of the spatial filters with multi-channel mixture signals [11]. Subsequently, the Transform-Average-Concatenate (TAC) mechanism is introduced in FasNet to avoid selecting a reference microphone [12]. However, the location information of the sound source has not been utilized to assist in the separation. In [13], an all-deep-learning MVDR framework is proposed to estimate every component in the calculation of an MVDR filter. Liu et al. utilizes the inter-channel phase difference (IPD) to convert the DOA into direction features to direct the separation network [14]. This kind of method requires prior knowledge of the DOA of the sound source, and its performance significantly degrades when there is a large error in DOA estimation. Some neural beamformer methods do not need to know the sources’ DOA, but they require the sources to remain spatially static [15],[16].
@@
Another kind of method to utilize the spatial information is training the multi-channel separation networks end to end with feature extraction across the channel dimension. Lee et al. extends the Conv-Tasnet [5] and extracts the inter-channel features with 2-D TCN layers [17]. A triple-path method utilizing Squeeze-and-Excitation mechanism [18] across the channel dimension to learn the spatial features is introduced in [19]. In [20], the dual-path transformer method [21] is further extended to a triple-path attentive recurrent network in which the ARN module consisting of a RNN block, an attention block and a feedforward block is applied on the inter-channel, intra-chunk and inter-chunk dimension. TF-GridNet is proposed using 2-D convolutional layers to extract spatial information, bidirectional Long Short-Term Memory networks (BLSTM) to extract intra-frame and sub-band information and multi-head self-attention to extract full-band information in [22]. SpatialNet is proposed in [23] which utilizes the convolutional layers to extract spatial information followed by cross-band blocks and narrow-band blocks operated on T-F dimension and outperforms TF-GridNet with fewer parameters. The state-space model [24] is further introduced to enhance training and inference speed in [25]. However, such methods only exploit spatial information implicitly by leveraging the multi-channel information and do not effectively utilize the DOA information of the sound source. Besides, these methods are only validated under stationary sound source scenarios.
@@
In fact, humans can effectively utilize spatial information in complex environments to simultaneously localize and separate sound sources. The localization results can aid in sound source separation, and the separation results can also assist the localization. Localization and separation are mutually supportive processes. Some studies have explored how to leverage both localization and separation simultaneously. In [26], the spatial area is initially divided into several regions, the sound of sources in each region is separated, and the boundaries of these regions are iteratively refined through binary search based on the iterative separation results. However, this method relies on the assumption that the sound source is stationary throughout the temporal duration of the input audio segment. Some methods implement source separation through predefined spatial partitions (e.g., distance-based separation) to separate sources within designated zones [27],[28],[29],[30]. Such methods exhibit constrained precision in spatial information utilization. Furthermore, when moving sources traverse inter-region boundaries, these algorithms exhibit separation failures due to spatial ambiguity. Taherian et al. perform localization based on multi-channel separation and used the localization results to determine the number of sound sources in the current frame, making the subsequent continuous separation more precise [31]. But this method does not incorporate the localization results into the separation process. LaSNet is proposed in [32], which integrates localization and separation within a single network, using loss functions for both tasks to jointly optimize the network. This approach considers localization and separation as a multi-task-learning problem, leveraging only the shared information between the two tasks in the hidden layer space, without clearly defining the mechanism by which the two tasks assist each other. This method also requires the sources to remain stationary.
@@
Most of the sound separation methods for handling moving sound sources mostly rely on dynamically estimating the location of the sound source. Some methods estimate the location of the sources with the steered response power (SRP) methods on each frame first and track the sources based on the estimated locations. The dynamic locations are utilized as steer vectors and the sound is separated based on them [33],[34]. The localization results may deteriorate under reverberant environments, which can significantly impact the separation performance. Besides, these methods do not leverage the relevant information between localization and separation. Munakata et al. jointly perform the sound separation and localization with an inference model with time-varying spatial covariance matrices (SCMs) and introduce the temporal smoothness of SCMs to the inference model [35]. However, this method relies on assumptions about the prior distribution of the data and the sound source separation and localization do not benefit from each other.
Inspired by coarse-to-fine iterative refinement strategies in some monaural separation approaches [36],[37], in this paper, a method utilizing Mutual Facilitation mechanism between sound source Tracking and Separation (MFTS) is proposed. The proposed method consists of three stages: initial tracking, mutual facilitation and neural beamforming. In the initial tracking stage, we propose an envelope-based tracking method. We first estimate the envelope signal of each sound source, which can be viewed as simplified sound separation. These estimated envelope signals then assist the sound source tracking. The initial tracking trajectory can serve as a starting clue for mutual facilitation. During the mutual facilitation stage, the separated multi-channel signal for a source is first derived from the estimated coarse trajectory, which in turn generates a refined trajectory; these two steps iteratively refine each other through alternating updates. Finally, the neural beamforming stage utilizing the refined multi-channel separation result and trajectory achieves enhanced accuracy in the single-channel source separation.
@@
The rest of this paper is organized as follows. Section II introduces the problem formulation of the target sound extraction. Section III presents the proposed method, including the tracking and sound extraction modules, and shows how they collaborate. Section IV verifies the effectiveness of the method through simulation experiments. Finally, our conclusion is presented in Section V.
@@ II Universal Sound Separation
$\begin{split}y_{m}(n)&=\sum_{c=0}^{C}x_{cm}(n)=\sum_{c=0}^{C}s_{c}(n)*h_{cm}(n),\end{split}$ (1)
where $y_{m}(n)$ is the mixture signal of microphone $m$ at time $n$, $C$ is the number of the sources, $x_{cm}(n)$ is the signal at the microphone $m$ from the $c$-th source, $s_{c}(n)$ is the signal of source $c$ and $h_{cm}(n)$ is the time-varying room impulse responses (RIR) between source $c$ and microphone $m$. The goal of universal sound separation is to separate the signal $\mathbf{x}_{cm}$ of sound source $c$ at reference channel $m$. The location of the sound sources is a clue that can effectively aid in sound separation. By providing the location or DOA of source $c$ as the additional input, the sound of the target source $c$ at microphone $m$ can be extracted.
The choice of reference microphone can also affect the performance of sound source separation. To eliminate the impact of reference microphone selection on separation performance for sources at different locations, we use First-Order Ambisonics (FOA) signals and the zeroth-order signal is selected as the separation target. The zeroth-order ambisonic signal is equivalent to the signal recorded by an omnidirectional microphone placed at the center of a spherical microphone array, thereby neutralizing the performance variations due to different source positions and reference microphone selections. Concurrently, by transforming the microphone array signals into the spherical harmonic domain, the algorithm achieves the decoupling of the spatial and frequency information and provides an elegant framework for spatial character exploitation [38].
Figure 1: The overall system of proposed method.
Figure 2: The schematic diagram of each stage. (a) Stage 1: the initial tracking module. (b) Stage 2: the mutual facilitation module. (c) Stage 3: the neural beamforming module.
@@ III Method
The overall schematic diagram of our proposed method is shown in Fig. 1. The proposed method can be divided into three stages. In the first stage, we perform initial tracking of all the sound sources based on the estimated envelope signal of each source. These initial trajectories might not be highly precise. In the second stage, a mutual facilitation process for source tracking and target sound extraction is performed. Specifically, the initial trajectory is used as a starting clue and input together with the mixture signals into the target source extraction network to extract the multi-channel sound corresponding to the trajectory. The separation result is subsequently input to precise tracking network and a precise tracking result based on the cleaner signal is generated. After that, this precise trajectory is then used as a clue to derive a refined multi-channel separation result based on which a refined trajectory can be then derived. In the third stage, a neural-beamformer utilizes the refined separation and tracking results to produce the single-channel sound separation signal. Fig. 2 shows the schematic diagram of each stage. The training mechanisms for the above networks are described as follows.
Figure 3: The structure of modified SpatialNet for envelope estimation.
@@ III-A Initial tracking
The initial tracking module takes mixture signals as input and outputs the trajectories of the sources. We first estimate the envelope signal of each sound source in the mixture. This envelope signal is then input to the tracking network along with the audio mixture signal. Concurrently, the tracking network’s output is multiplied by the envelope signal to compute the loss.
III-A1 Envelope estimator
@@
The groundtruth envelope for the signal of the $c$-th source is derived through a max-pooling layer applied to the absolute value of the single-channel zeroth-order FOA signal. The max-pooling layer operates by extracting frame-wise peak values. The envelope of the $c$-th source can be denoted as $\mathbf{\hat{E}}^{trg}_{c}\in\mathbb{R}^{T\times 1}$, where $T$ is the number of frames. The full-length envelope signal of the $c$-th source , which temporally aligns with the audio mixture, is obtained by linearly interpolating $\mathbf{\hat{E}}^{trg}_{c}$, which is denoted as $\mathbf{E}^{trg}_{c}\in\mathbb{R}^{N\times 1}$, where $N$ is the number of temporal samples of input audio signal. $N=fs\cdot L$, where $f_{s}$ is the sampling rate (Hz) and $L$ is the duration of the signal (second). The network used for estimating the envelopes in this study is SpatialNet [23], which is a network developed for sound separation. SpatialNet accepts multi-channel mixed signals as input and outputs single-channel separated signal for all sound sources. We select this network as the task of estimating the envelope of each sound source from the audio mixture can be considered as a simplified version of sound separation. The structure of SpatialNet for envelope estimation is shown in Fig. 3. The mixed signal is transformed to T-F domain by Short Time Fourier Transform (STFT), which is denoted as $\mathbf{Y}\in\mathbb{R}^{F\times T\times 2M}$, where $M=4$ is the number of orders of the FOA and $T$ is the number of frames and $F$ is the number of frequency points. The T-F domain representation first undergoes a 1-D convolutional layer to extract inter-channel features. It is subsequently processed by the Cross-band block to capture frequency-band interaction features correlated with source locations. Within the Cross-band block, the input features sequentially pass through one frequency-domain group convolutional layer, one frequency-domain group of linear layers, and another frequency-domain group convolutional layer. The features are then passed through the Narrow-band block to extract intra-band temporal information. Following the method in [25], we employ Mamba to replace the original multi-head self-attention layer and the temporal convolutional layer in SpatialNet, thereby reducing computational load. Specifically, the features sequentially undergo two Mamba layers bracketed by linear layers. Finally, the features are processed through a linear projection layer, inverse STFT (iSTFT), and a max-pooling layer to obtain the estimated envelopes, $\mathbf{\hat{E}}^{est}\in\mathbb{R}^{C\times T\times 1}$, of $C$ separated sources. The loss function for training the envelope estimation network is Normalized Mean Square Error (NMSE), which is denoted as:
$\begin{split}\mathcal{L}_{env}&=\text{NMSELoss}(\mathbf{\hat{E}}^{est},\mathbf{\hat{E}}^{trg})\\ &=\frac{1}{C}\sum_{c=0}^{C}(10\log_{10}\frac{\sum_{t=0}^{T}(\hat{E}^{est}_{c}(t)-\hat{E}^{trg}_{c}(t))^{2}}{\sum_{t=0}^{T}(\hat{E}^{trg}_{c}(t))^{2}}),\end{split}$ (2)
@@
where $\mathbf{\hat{E}}^{trg}\in\mathbb{R}^{C\times T\times 1}$ denotes the groundtruch envelopes for all the sources. Besides, utterance-level permutation invariant training (uPIT) is applied during training. The estimated envelope signals are also obtained by linearly interpolating $\mathbf{\hat{E}}^{est}$ along the temporal dimension , which is denoted as $\mathbf{E}^{est}\in\mathbb{R}^{C\times N\times 1}$.
III-A2 Initial tracking